rebeccaxhawaii (rebeccaxhawaii):

how to make cereal 1 carton of cereal 1 carton of milk 1 bowl 1 spoon first put the cereal in the bowl then put the milk inside finally put the spoon in the bowl I have only tried this with Froot loops so I don't know if it works with other cereal

rebeccaxhawaii (rebeccaxhawaii):

math question is Let’s use this function to represent the lizard population. L(t) = 14(2)^t The lizard population starts at 14 and doubles every year. However, it is now necessary to find the population each month.

OpenStudy (anonymous):

I don't know the meaning ^

OpenStudy (anonymous):

how many months in a year?

rebeccaxhawaii (rebeccaxhawaii):

12

OpenStudy (anonymous):

ok you know it, so lets cut to the chase

OpenStudy (anonymous):

in one month it will be \[\huge 15\times 2^{\frac{t}{12}}\]

OpenStudy (anonymous):

oops that was suppose to be a 14, not a 15

rebeccaxhawaii (rebeccaxhawaii):

its good i can pretend but oki

OpenStudy (anonymous):

if you want a formula for each month, then do \[14\times \left(2^{\frac{1}{12}}\right)^t\]

OpenStudy (anonymous):

you can find \(2^{\frac{1}{12}}\) with a calculator

rebeccaxhawaii (rebeccaxhawaii):

wouldnt it be

rebeccaxhawaii (rebeccaxhawaii):